# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ Abstract
1 Introduction
Quantum Key Distribution (QKD) [1, 2, 3, 4, 5] allows two parties to generate a secret shared key whose unconditional security is guaranteed by the principles of quantum mechanics. In recent decades, many security proofs for QKD have been devised (e.g., [6, 7, 8, 9, 10, 11]), proving security under specific theoretical assumptions.
QKD schemes have not only been described theoretically, but also implemented in practice, including in laboratory experiments and commercial products [12, 13, 14, 5, 15].
Unfortunately, those real-world implementations of QKD inevitably deviate from the idealized models assumed in security proofs, leading to possibly exploitable loopholes. As a simple example, while most QKD schemes are based on qubits, they are implemented by sending and receiving photons, which actually reside in a quantum space of higher dimension. Such mismatches between theoretical models and practical implementations of QKD have given rise to various attacks that effectively compromise the security of the secret key [16, 17, 18, 19, 20, 21, 22, 23, 24]; these attacks highlight the need for a deeper understanding of security in practical QKD systems and better ways to analyze their security.
In contrast to the analysis of QKD imperfections, the security analysis and certification of practical implementations of communication systems have been thoroughly researched in the classical field of cybersecurity, resulting in well-established methods for such analyses [25, 26]. The first step in such an analysis is to review the interfaces of the system with potential adversaries, interfaces named the attack surface. The next step (called “vulnerability research”) is to search for vulnerabilities, which are unwanted behaviors of the system, some related to the protocol and some to the environment, the latter called side channels. The final step is to build an exploit: an attack that uses the detected vulnerabilities.
@@ 1 Introduction
Then, we define the first vulnerability research method for QKD implementations, which we name Quantum Fuzzing (see Section 4). This method is analogous to the notion of classical fuzzing, which “experiments” with unusual inputs to an implementation in order to expose vulnerabilities. Adapting this concept to the quantum world, we enable QKD system designers and attackers to discover vulnerabilities in a QKD system, even if their knowledge of its inner workings is limited.
@@
We next define a general family of attacks named Reversed-Space Attacks (Section 5), which is a generic exploit method for discovered vulnerabilities. This method, initially introduced in the preliminary work of [27, 28], begins by finding the reversed space of the QKD implementation, which is derived by reversing states that are measured by the QKD device “back in time”. This space is essentially the receiving party’s attack surface, because it exactly describes the quantum space affecting their measurements. Then, the method uses the reversed space to construct a successful attack (see Appendix B for a detailed example). We note that an additional important example of a reversed-space attack is the “fixed-apparatus attack”, which appeared in [24].
@@
In contrast to attacks that only rely on the transmitted and received quantum states, there are also attacks that utilize unmeasured side channels; such attacks are named “side-channel attacks”. Side-channel attacks on quantum devices have appeared in the literature [20, 21, 29, 30]; however, previous definitions of this term were too general, because all attacks on QKD systems rely on physical faults. We thus give new and meaningful definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks (see Section 6).
@@
To demonstrate how the methods we developed can help devise and analyze attacks on practical QKD implementations, we show that the Bright Illumination attack [31, 21, 32] is, in fact, a delicately crafted exploitation of the tools defined in this paper (see Section 7). Finally, we use our tools to classify well-known attacks on practical QKD implementations (see Section 8).
@@
Initial versions of (subsets of) the results presented in this paper previously appeared in [27] (see also extended results in [28]) and [32].
1.1 Other Related Work
@@
Our work focuses on implementations of QKD schemes using photons. This is not a limitation, since all commercial implementations and almost all experimental implementations of QKD are photon-based [12, 13, 14, 5, 15, 33, 34, 35, 36, 37].
@@
Many attacks on QKD implementations have appeared in the literature (see, e.g., [16, 17, 18, 21, 22, 23, 24]; see Appendix G for more details). Previous works have also classified certain similar attacks together as “attack families”, which are sets of attacks that share common principles; two examples include the “faked states” attack family [19, 20] and the “Detector Efficiency Mismatch” attack family [38].
@@
A thorough survey of attacks on QKD implementations was given in [30]. In addition, the term “vulnerabilities” in reference to QKD imperfections was also used in the review [39].
@@
Finally, in this paper, we use classical cybersecurity terminology that has been commonplace in the last few decades. See [25] for an educational handbook and the NIST Special Publication 800-160 documents [26] for cybersecurity guidelines of the US federal government: both provide documentation of that terminology, as well as a broad introduction to classical cybersecurity as a whole.
2 Preliminaries
2.1 The (Ideal) BB84 Protocol
In the BB84 QKD protocol [1], whose two legitimate participants are typically named Alice and Bob, Alice sends to Bob $N$ qubits, each of them randomly chosen from the set of quantum states $\{|{0}\rangle,|{1}\rangle,|{+}\rangle\triangleq\frac{|{0}\rangle+|{1}\rangle}{\sqrt{2}},|{-}\rangle\triangleq\frac{|{0}\rangle-|{1}\rangle}{\sqrt{2}}\}$, and Bob measures each of them either in the computational basis $\{|{0}\rangle,|{1}\rangle\}$ or in the Hadamard basis $\{|{+}\rangle,|{-}\rangle\}$, chosen randomly. Thereafter, Alice and Bob post-process the results by using an authenticated classical channel. If Alice and Bob used matching bases, they now share a bit (unless there is some noise or eavesdropping); if they use mismatching bases, Bob’s results are random. Alice and Bob reveal their basis choices and discard the bits for which they used mismatching bases. After that, they publicly reveal a random subset of their bits in order to estimate the error rate (then discarding those exposed test bits), aborting the protocol if the error rate is too high; and they perform error correction and privacy amplification processes, obtaining a completely secret, identical final key.
2.2 Realistic QKD Implementations and Fock Space Notation
@@
While the ideal BB84 (and other QKD protocols) communicates using abstract qubits, any experimental implementation must rely on physical, easy-to-transfer quantum carriers, which encode qubits. In most existing implementations, photons serve as such carriers. A very well-known example (e.g., [12]) is the BB84 implementation that uses two polarization modes to encode a qubit: namely, the polarizations $\leftrightarrow$ and $\updownarrow$ signify $|{0}\rangle$ and $|{1}\rangle$ respectively, and the orthogonal diagonal photon polarizations encode $|{+}\rangle$ and $|{-}\rangle$.
@@
Using this notation, the above polarization-based implementation can be expressed using two modes, i.e., one mode to signify photons in the $\leftrightarrow$ polarization and another mode for $\updownarrow$. The state $|{0}\rangle=|{0,1}\rangle^{\!\text{\tiny F}}$ describes a single photon with horizontal polarization, and $|{1}\rangle=|{1,0}\rangle^{\!\text{\tiny F}}$ describes a single photon with vertical polarization. Similarly, in implementations that encode a qubit onto the time at which a photon arrives, we can define two distinct time bins: $t_{0}$ and $t_{1}$. In this case, $|{0}\rangle=|{0,1}\rangle^{\!\text{\tiny F}}_{t_{1},t_{0}}$ is a photon arriving at the first time-bin, and $|{1}\rangle=|{1,0}\rangle^{\!\text{\tiny F}}_{t_{1},t_{0}}$ is a single photon arriving at the second time-bin. We will use $|{V}\rangle=|{0,0}\rangle^{\!\text{\tiny F}}$ to denote the vacuum state — a state with zero photons in all modes. Note that since photons are indistinguishable and have integer spin, the state remains identical when photons “switch places”, and there is no upper bound on the number of photons in each mode.
@@
In addition, one can measure other specific properties of the state using linear optical tools, such as beam splitters, phase shifters, and mirrors (see [40]). Details on the applications of these tools on photonic quantum states can be found in Appendix A.
3 Classical and Quantum Cybersecurity
@@ 3.1 Classical Cybersecurity
    An attack surface is “the set of points on the boundary of a system, a system element, or an environment where an attacker can try to enter, cause an effect on, or extract data from”. [26]
@@
    A vulnerability is “a flaw in a system’s security that can lead to an attacker utilizing the system in a manner other than that which the designer intended”. [25]
@@
    An example of a well-known vulnerability is a “Buffer Overflow” (BOF) (see, e.g., [41]), where a program writes data to an array of bytes (a buffer) that is shorter than the data length. This causes the data to overflow from the confines of the dedicated buffer to the next memory cells, affecting other data in the computer’s memory. In the context of our blog website, such a vulnerability may occur in the code handling the “post comment” request if a user can send, for example, an arbitrarily long email address while the server allocates a fixed-size buffer to store it.
@@
    An exploit is a “tool, set of instructions, or code that is used to take advantage of a vulnerability”. [25].
The above concepts are highly beneficial for analyzing attacks on computing systems because of their ability to decompose attacks into separate stages and components. Once each component is considered separately, researchers can better understand how this component behaves, find more instances of it, and build tools to defend against it. For example, researching vulnerabilities has given rise to definitions of common types of vulnerabilities [41], methods for discovering vulnerabilities in code [42], and coding practices that minimize vulnerabilities [43]. Researching attack surfaces has given rise to attack surface characterization methods [44], as well as tools for minimizing attack surfaces (such as firewalls). Researching the concept of exploits has given rise to generic exploit methods [45], as well as defense mechanisms that prevent exploits [46, 47].
3.2 Cybersecurity of QKD systems
We can now analyze the security of QKD systems through the lens of the (classical) cybersecurity analysis framework depicted above, extending the classical framework to QKD systems. This allows us to decompose complete attacks into their constituent components and analyze each one in detail, as is standard practice in classical cybersecurity.
3.2.1 Vulnerabilities
@@
A common form of vulnerability in Alice’s device is that Alice sometimes sends different states than those she intended. A well-known example of such a vulnerability can be seen in the Photon-Number Splitting (PNS) attack [16, 48] (which showed all QKD experiments done until around year 2000 to be insecure). The PNS vulnerability is based on a fault in Alice’s transmitter that sometimes sends two-photon pulses (instead of a single photon) in one of the four allowed configurations.
@@
Suppose Alice sends a perfect qubit encoded into the polarization of a single photon, and yet Bob uses a detector that cannot distinguish a single photon from a pair of photons (commonly called a “threshold detector”). Thus, while Alice only sends states from the Hilbert space $\operatorname{span}\{|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}}\}$, if the state $|{2,0}\rangle^{\!\text{\tiny F}}$ arrives at Bob’s detector, the detector cannot distinguish it from the state $|{1,0}\rangle^{\!\text{\tiny F}}$. Thus, Eve can attack a larger space than is possible in the ideal case. [^2] There are known ways to mitigate this vulnerability, such as through the squashing model [49].
@@
Suppose Alice sends a perfect qubit encoded to the polarization of a single photon, which arrives at a specific time $t$, and yet Bob cannot exactly distinguish when a photon arrives. Namely, for some delay $\delta>0$, he treats late photons that arrive at $t+\delta$ the same as photons that arrive correctly at time $t$. Modeling this example using Fock states, with four modes describing “horizontal at tt”, “vertical at tt”, “horizontal at $t+\delta$” and “vertical at $t+\delta$”, Alice’s qubit is encoded using the span of $\{|{1,0,0,0}\rangle^{\!\text{\tiny F}},|{0,1,0,0}\rangle^{\!\text{\tiny F}}\}$, but Bob interprets $|{0,0,1,0}\rangle^{\!\text{\tiny F}}$ the same as $|{1,0,0,0}\rangle^{\!\text{\tiny F}}$ and interprets $|{0,0,0,1}\rangle^{\!\text{\tiny F}}$ the same as $|{0,1,0,0}\rangle^{\!\text{\tiny F}}$.

### v3

@@ Abstract.
Quantum Key Distribution, Cybersecurity, Implementation Attacks
$c$s: Hardware Quantum communication and cryptographyccs: Security and privacy Cryptanalysis and other attacksccs: Security and privacy Hardware attacks and countermeasuresccs: Theory of computation Quantum information theoryccs: Security and privacy Key managementccs: Security and privacy Hardware-based security protocols
1. Introduction
Quantum Key Distribution (QKD) (BB84, ; B92, ; E91, ; COW-QKD, ; Xu2020, ) allows two parties to generate a secret shared key whose unconditional security is guaranteed by the principles of quantum mechanics. In recent decades, many security proofs for QKD have been devised (e.g., (MayersBB84Sec, ; ShorBB84Sec, ; BBBMR00_conf, ; BBBMR06, ; RennerBB84SecCompose, ; BenOrBB84SecCompose, )), proving security under specific theoretical assumptions.
QKD schemes have not only been described theoretically, but also implemented in practice, including in laboratory experiments and commercial products (bennett_experimental_1992, ; muller_plug_1997, ; Stucki_2009, ; Xu2020, ; Peev2009, ).
Unfortunately, those real-world implementations of QKD inevitably deviate from the idealized models assumed in security proofs, leading to possibly exploitable loopholes. As a simple example, while most QKD schemes are based on qubits, they are implemented by sending and receiving photons, which actually reside in a quantum space of higher dimension. Such mismatches between theoretical models and practical implementations of QKD have given rise to various attacks that effectively compromise the security of the secret key (PNS, ; vakhitov_large_2001, ; makarov_faked_2005, ; gisin_trojan-horse_2006, ; qi_time-shift_2006, ; makarov_faked_2008, ; lydersen_blinding_commercial_2010, ; boyer_fixedapp_2014, ; Pang2020, ); these attacks highlight the need for a deeper understanding of security in practical QKD systems and better ways to analyze their security.
In contrast to the analysis of QKD imperfections, the security analysis and certification of practical implementations of communication systems have been thoroughly researched in the classical field of cybersecurity, resulting in well-established methods for such analyses (shellcoders_handbook, ; Ross2021_SP800, ). The first step in such an analysis is to review the interfaces of the system with potential adversaries, interfaces named the attack surface. The next step (called “vulnerability research”) is to search for vulnerabilities, which are unwanted behaviors of the system, some related to the protocol the system implements and some to the environment, the latter called side channels. The final step is to build an exploit: an attack that uses the detected vulnerabilities.
@@ 1. Introduction
Then, we define the first vulnerability research method for QKD implementations, which we name Quantum Fuzzing (see Section 4). This method is analogous to the notion of classical fuzzing (fuzzing, ; fuzzing_review_1, ; fuzzing_review_2, ), which “experiments” with unusual inputs to an implementation in order to expose vulnerabilities. Adapting this concept to the quantum world, we enable QKD system designers and attackers to discover vulnerabilities in a QKD system, even if their knowledge of its inner workings is limited.
@@
We next define a general family of attacks named Reversed-Space Attacks (Section 5), which is a generic exploit method for discovered vulnerabilities. This method, initially introduced in the preliminary work of (gelles_security_2012, ; gelles_reversed_2016, ), begins by finding the reversed space of the QKD implementation, which is derived by reversing states that are measured by the QKD device “back in time”. This space is essentially the receiving party’s attack surface, because it exactly describes the quantum space affecting their measurements. Then, the method uses the reversed space to construct a successful attack (see Appendix B for a detailed example). We note that an additional important example of a reversed-space attack is the “fixed-apparatus attack”, which appeared in (boyer_fixedapp_2014, ).
@@
In contrast to attacks that only rely on the transmitted and received quantum states, there are also attacks that utilize unmeasured side channels; such attacks are named “side-channel attacks”. Side-channel attacks on quantum devices have been discussed in the literature (makarov_faked_2008, ; lydersen_blinding_commercial_2010, ; after_gate, ; bsi, ) as attacks that rely on physical faults; however, these previous definitions were too general, because all attacks on QKD systems rely on physical faults. We thus give new and meaningful definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks (Section 6).
@@
To demonstrate how the methods we developed can help devise and analyze attacks on practical QKD implementations, we show that the Bright Illumination attack (makarov_blinding_first_2009, ; lydersen_blinding_commercial_2010, ; liss_practice_2020, ) is, in fact, a delicately crafted exploitation of the tools defined in this paper (Section 7). Finally, we use our tools to classify well-known attacks on practical QKD implementations (Section 8).
@@
Initial versions of (subsets of) the results presented in this paper previously appeared in (gelles_security_2012, ) (see also extended results in (gelles_reversed_2016, )) and (liss_practice_2020, ).
1.1. Other Related Work
@@
Our work focuses on implementations of QKD schemes using photons. This is not a limitation, since all commercial implementations and almost all experimental implementations of QKD are photon-based (bennett_experimental_1992, ; muller_plug_1997, ; Stucki_2009, ; Xu2020, ; Peev2009, ; Yuan2018, ; Zhang2020, ; Chen2021, ; Kwek2021, ; BassoBasset2021, ).
@@
Many attacks on QKD implementations have appeared in the literature (see, e.g., (PNS, ; vakhitov_large_2001, ; gisin_trojan-horse_2006, ; lydersen_blinding_commercial_2010, ; Pang2020, ; qi_time-shift_2006, ; boyer_fixedapp_2014, ); see Appendix E for more details). Previous works have also classified certain similar attacks together as “attack families”, which are sets of attacks that share common principles; two examples include the “faked states” attack family (makarov_faked_2005, ; makarov_faked_2008, ) and the “Detector Efficiency Mismatch” attack family (makarov_effects_2006, ).
@@
A thorough survey of attacks on QKD implementations was given in (bsi, ). In addition, the term “vulnerabilities” in reference to QKD imperfections was also used in the review (Zapatero2025, ).
@@
Finally, in this paper, we use classical cybersecurity terminology that has been commonplace in the last few decades. See (shellcoders_handbook, ) for an educational handbook and the NIST Special Publication 800-160 documents (Ross2021_SP800, ) for cybersecurity guidelines of the US federal government: both provide documentation of that terminology, as well as a broad introduction to classical cybersecurity as a whole.
2. Preliminaries
2.1. The (Ideal) BB84 Protocol
The first and most prominent QKD protocol is BB84, invented by Bennett and Brassard (BB84, ). It allows two legitimate participants, typically named Alice and Bob, to generate a secret shared key.
The protocol begins with a quantum communication phase, where Alice sends $N$ transmissions to Bob. In each transmission, Alice randomly chooses a quantum state out of the set $\{|{0}\rangle,|{1}\rangle,|{+}\rangle\triangleq\frac{|{0}\rangle+|{1}\rangle}{\sqrt{2}},|{-}\rangle\triangleq\frac{|{0}\rangle-|{1}\rangle}{\sqrt{2}}\}$, and Bob randomly chooses whether to measure in the computational basis $\mathrm{B}_{\mathrm{C}}=\{|{0}\rangle,|{1}\rangle\}$ or the Hadamard basis $\mathrm{B}_{\mathrm{H}}=\{|{+}\rangle,|{-}\rangle\}$. Then, Alice transmits her chosen state, and Bob measures it in his chosen basis.
After all transmissions have been completed, Alice and Bob continue the protocol using a classical authenticated communication channel (to which where Eve can listen, but not interfere). First, Alice reveals whether each of her transmissions was made in the computational or the Hadamard basis, and Bob reveals which basis he used for measuring the transmission. Alice and Bob then discard all transmissions where they used mismatching bases (since the results carry no information) and retain a set of measurement results that should be equal if no eavesdropping or noise occurred.
Thereafter, Alice and Bob choose a random subset of their measurement results and publish it over the classical channel, estimating the amount of mismatches in their shared string. The fraction of errors in transmission (that is, the number of transmissions that had errors, divided by the total number of transmissions) is called the quantum bit error rate (QBER). If the QBER is above a certain threshold, an eavesdropper has completely compromised the security of the transmissions, and the protocol is aborted; otherwise, Alice and Bob continue the protocol with the non-published secret bits, defined as the raw key. Since the raw key can still include mismatches between Alice and Bob, and since Eve can still have little information about it, Alice and Bob perform classical phases of error correction and privacy amplification to transform the raw key into a shorter string (the final key) which is completely equal for Alice and Bob and (up to a negligible probability) completely secret from Eve.
2.2. Realistic QKD Implementations and Fock Space Notation
@@
While the ideal BB84 (and other QKD protocols) communicates using abstract qubits, any experimental implementation must rely on physical, easy-to-transfer quantum carriers, which encode qubits. In most existing implementations, photons serve as such carriers. A very well-known example (e.g., (bennett_experimental_1992, )) is the BB84 implementation that uses two polarization modes to encode a qubit: namely, the polarizations $\leftrightarrow$ and $\updownarrow$ signify $|{0}\rangle$ and $|{1}\rangle$ respectively, and the orthogonal diagonal photon polarizations encode $|{+}\rangle$ and $|{-}\rangle$.
@@
Using this notation, the above polarization-based implementation can be expressed using two modes: one mode signifying photons in the horizontal polarization (denoted $\leftrightarrow$ or $\mathrm{H}$) and another mode signifying photons in the vertical polarization (denoted $\updownarrow$ or $\mathrm{V}$). The state $|{0}\rangle=|{0{,}1}\rangle^{\!\text{\tiny F}}_{\mathrm{V},\mathrm{H}}$ describes a single photon with horizontal polarization, and $|{1}\rangle=|{1{,}0}\rangle^{\!\text{\tiny F}}_{\mathrm{V},\mathrm{H}}$ describes a single photon with vertical polarization. Similarly, in implementations that encode a qubit onto the time at which a photon arrives, we can define two distinct time bins: $t_{0}$ and $t_{1}$. In this case, $|{0}\rangle=|{0{,}1}\rangle^{\!\text{\tiny F}}_{t_{1},t_{0}}$ is a photon arriving at the first time-bin, and $|{1}\rangle=|{1{,}0}\rangle^{\!\text{\tiny F}}_{t_{1},t_{0}}$ is a single photon arriving at the second time-bin. We will use $|{V}\rangle=|{0{,}0}\rangle^{\!\text{\tiny F}}$ to denote the vacuum state — a state with zero photons in all modes. Note that since photons are indistinguishable and have integer spin, the state remains identical when photons “switch places”, and there is no upper bound on the number of photons in each mode.
@@
In addition, one can measure other specific properties of the state using linear optical tools, such as beam splitters, phase shifters, and mirrors (see (RZBB94, )). Details on the applications of these tools on photonic quantum states can be found in Appendix A.
3. Classical and Quantum Cybersecurity
@@ 3.1. Classical Cybersecurity
    An attack surface is “the set of points on the boundary of a system, a system element, or an environment where an attacker can try to enter, cause an effect on, or extract data from”. (Ross2021_SP800, )
@@
    A vulnerability is “a flaw in a system’s security that can lead to an attacker utilizing the system in a manner other than that which the designer intended”. (shellcoders_handbook, )
@@
    An example of a well-known vulnerability is a “Buffer Overflow” (BOF) (see, e.g., (aleph1996smashing, )), where a program writes data to an array of bytes (a buffer) that is shorter than the data length. This causes the data to overflow from the confines of the dedicated buffer to the next memory cells, affecting other data in the computer’s memory. In the context of our blog website, such a vulnerability may occur in the code handling the “post comment” request if a user can send, for example, an arbitrarily long email address while the server allocates a fixed-size buffer to store it.
@@
    An exploit is a “tool, set of instructions, or code that is used to take advantage of a vulnerability”. (shellcoders_handbook, ).
The above concepts are highly beneficial for analyzing attacks on computing systems because of their ability to decompose attacks into separate stages and components. Once each component is considered separately, researchers can better understand how this component behaves, find more instances of it, and build tools to defend against it. For example, researching vulnerabilities has given rise to definitions of common types of vulnerabilities (aleph1996smashing, ), methods for discovering vulnerabilities in code (fuzzing, ), and coding practices that minimize vulnerabilities (seacord2013secure, ). Researching attack surfaces has given rise to attack surface characterization methods (Manadhata_metric_2011, ), as well as tools for minimizing attack surfaces (such as firewalls). Researching the concept of exploits has given rise to generic exploit methods (shacham2007geometry, ), as well as defense mechanisms that prevent exploits (shacham2004effective, ; abadi2005cfi, ).
3.2. Cybersecurity of QKD systems
We can now analyze the security of QKD systems through the lens of the classical cybersecurity analysis framework depicted above, extending the framework to QKD systems. This allows us to decompose complete attacks into their constituent components and analyze each one in detail, as is standard practice in classical cybersecurity.
3.2.1. Vulnerabilities
@@
A common form of vulnerability in Alice’s device is that Alice sometimes sends different states than those she intended. A well-known example of such a vulnerability can be seen in the Photon-Number Splitting (PNS) attack (PNS, ; PNS_eurocrypt, ) (which showed all QKD experiments done until around year 2000 to be insecure). The PNS vulnerability is based on a fault in Alice’s transmitter that sometimes sends two-photon pulses (instead of a single photon) in one of the four allowed configurations.
@@
Suppose Alice sends a perfect qubit encoded into the polarization of a single photon, and yet Bob uses a detector that cannot distinguish a single photon from a pair of photons (commonly called a “threshold detector”). Thus, while Alice only sends states from the Hilbert space $\operatorname{span}\{|{0{,}1}\rangle^{\!\text{\tiny F}},|{1{,}0}\rangle^{\!\text{\tiny F}}\}$, if the state $|{0{,}2}\rangle^{\!\text{\tiny F}}$ arrives at Bob’s detector, the detector cannot distinguish it from the state $|{0{,}1}\rangle^{\!\text{\tiny F}}$. Thus, Eve can attack a larger space than is possible in the ideal case. [^2] There are known ways to mitigate this vulnerability, such as using the squashing model (squashing, ).
@@
Suppose Alice sends a perfect qubit encoded to the polarization of a single photon, which arrives at a specific time $t$, and yet Bob cannot exactly distinguish when a photon arrives. Namely, for some delay $\delta>0$, he treats late photons that arrive at $t+\delta$ the same as photons that arrive correctly at time $t$.
We model this example using Fock states, with four modes describing “horizontal at tt” ($\mathrm{H}_{t}$), “vertical at tt” ($\mathrm{V}_{t}$), “horizontal at $t+\delta$” ($\mathrm{H}_{t+\delta}$), and “vertical at $t+\delta$” ($\mathrm{V}_{t+\delta}$): the most general Fock state is denoted by $|{m_{3},m_{2},m_{1},m_{0}}\rangle^{\!\text{\tiny F}}_{\mathrm{V}_{t+\delta},\mathrm{H}_{t+\delta},\mathrm{V}_{t},\mathrm{H}_{t}}$. Alice’s qubit is encoded using the span of $\{|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}},|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}\}$, but Bob interprets $|{0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}$ the same as $|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}$ and interprets $|{1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}$ the same as $|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}$, because Bob can properly identify the polarization ($\mathrm{H}$ or $\mathrm{V}$) but cannot distinguish the time $t$ from $t+\delta$.